OpenStudy (anonymous):

A bike rental company charges $5 per hour per person for a bike. Each person riding a bike is required to wear a helmet. The company charges a flat rate of $8 per helmet. If two people rented bikes for three hours, how much did they pay? $13 $26 $40 $46

OpenStudy (anonymous):

$46

OpenStudy (anonymous):

8*2 = 16 (price of helmets) 5*3=15 ( price of one bike) 30=price of rental. So 30+16=46

OpenStudy (akashdeepdeb):

Two people require 2 helmets thus they pay = 16$ And they went for three hours thus they payed = 30$ Which is 46$ :) I hope this helped! :)

OpenStudy (anonymous):

@live10000000 This isnt a free answer site. If you give the answer show how you got it.

OpenStudy (anonymous):

live thank you but nathan explained it for me :)